Lisgar Ltd. is a Canadian company with a December 31 year end. On February 1, 2009, Lisgar acquires 15,000 shares of a Chinese corporation at a cost of 150 Yuan (Y, hereafter) per share. The total cost of the investment is Y2,250,000.

On June 30, 2009, the Chinese company declares and pays a dividend of Y15 per share.

When Lisgar closes its books on December 31, 2009, the fair value of the shares has increased to Y175 per share.

On August 1, 2010, the shares are sold for Y210 per share.

Relevant spot exchange rates for the Chinese Yuan are as follows:

February 1, 2009                                                                       Y1 = $0.145

June 30, 2009                                                                            Y1 = $0.150

December 31, 2009                                                                   Y1 = $0.170

August 1, 2010                                                                          Y1 = $0.165

Required:

A.        Provide dated journal entries to record the preceding information.

B.        Provide dated journal entries to record the preceding information assuming that Lisgar Ltd. classifies its investment as FVTOCI.
Answer #1

Feb 01, 2009 - Purchase of Investment

Number of share acquired = 15000

Cost of acquiring shares is 150Y

Total cost of investment = Number of shares*cost of each share= 15000*150 = 2,250,000Y

Date Journal entry Debit Credit
01 Feb 2009 Investment in share of Chinese Corporation A/c Dr. 3,26,250
To Cash paid 3,26,250
being cash paid the rate of Y1 = $0.145 so 2250000*0.145 = 326250

On June 30, 2009, Dividend declared and paid are the same date hence we are not accruing the dividends

Date Journal entry Debit Credit
30 June 2009 Cash received A/c Dr. 33,750
To Dividend income 33,750
being dividend received on 15000 shares at the rate of Y15 per share and the f/x from Y to $ being 0.150 = 15000*15*0.150 = 33,750

December 31, 2009 Increase in fair market value(FMV) of the shares we need to book unrealised gain

Date Journal entry Debit Credit
31 December 2009 Investment in share of Chinese Corporation A/c Dr. 63,750
To Unrealised gain 63,750
being unrealised gain of 25Y per share as the cost is 150Y and the FMV is 175 hence an unrealised gain of 25Y per shareand the f/x from Y to $ being 0.170 = 15000*25*0.170 = 63750

On August 1, 2010 - Sale of Investments of Chinese Corporation

Date Journal entry Debit Credit
01 August 2010 Cash received a/c Dr. 5,19,750
To Investment in share of Chinese Corporation A/c Dr. 3,26,250
To Realised gain on the sale of investment 148,500
To Realised F/X gain on the sale of investment 45,000
Being the realised gain on the sale of investment and realised f/x gain

Cash received = 15000 share * 210 sale price * 0.165 f/x rate for Y to $ 0.165 = 15000*210*0.165 = 5,19,750

Investment in share will be reversed at the original value recorded in the books 01 Feb 2019 which is 3,26,250

Realised gain = 15000*(sale price - cost) * f/x rate as on the sale of the investment = 15000*(210-150)*0.165 = 1,48,500

The balance amount is nothing but the gain arising due to F/X

Reversal of the unrealised gain as the investments are sold off

Date Journal Entry Debit Credit
01 August 2010 Unrealised gain A/c Dr 63,750
Investment in share of Chinese Corporation A/c 63,750
being reversal on account of sale of investment

Give a thumbs up if you are satisfied with the answer.